DeltaHcomb = -3910.3 kJ/mol for liquid toluene (C7H8, Mm = 92.70 g/mol) at 298 K. What temperature change would result at 298K from the combustion of 445.9 mg of liquid toluene in a constant pressure calorimeter with a heat capacity of 7.45 kJ·oC-1?

qcomb =  kJ

deltacap.gifT =

oC
Answer #1

mass of toluene = 445-9 mg = 445-9x103 gm molas mass of toluene (C7H8) = 92 g/mol a moles of toluene = 445091163 C = 4.85X10-